At a Friday night performance of a play, 120 adults and 75 student tickets were sold. The total receipts were $750. At the Saturday afternoon performance, 80 adult and 110 student tickets were sold for a total receipts of $620. Find the cost of a student ticket.
The answers are:
a)$1.50
b)$1.75
c)$2.00
d)2.25
I have an answer of $2.50.
120a+75s=$750
80a+110s=$620

Answer by checkley71(8403)

120a+75s=$750
80a+110s=$620 MULTIPLY THIS EQUATION BY -1.5 & ADD THEM.
-120A-165S=-930
120A+75S=750
---------------------
-90S=-180
S=-180/-90
S=$2 AMOUNT OF STUDENT TICKETS.